7 . You need to install a new cable to handle four years o f growth on a transmission

route. The traffic now fills one fiber, and traffic is doubling every three months.

How many fibers will you need in four years?

a. 4

b. 16

c. 128

d. 65,536

e. over 1 million

As in Problem 7, you need to install a new cable to handle four years of growth on a transmission route where traffic is doubling every three months. All the traffic now fits in a signal that requires one wavelength in a fiber that can handle 32 wavelengths with wavelength-division multiplexing. How many fibers will you need to carry that traffic if you fill each one with 32 wavelengths?

4

b. 16

c. 128

d. 2048

e. 65,536

9. Reality has set in, and you realize that traffic is doubling every year. How many

fibers would you need if each fiber carried only one signal and the first fiber was

already full?

a. 1

b. 4

c. 16

d. 128

e. 2048

1 0 . How many fibers would you need to

if each fiber could transmit signals at

handle the transmission load in Problem 9

32 wavelengths?

a. 1

b. 2

c. 4

d. 16

e. can’t tell from data given
Answer #1

7.
Since traffic doubles every 3 months hence in 4 years (16, 3-Months ) it will become 2^16.

And since we need 1 fibre for 1 signal it implies we need 2^16 = 65536 fibres. (D)

8. Now one fibre can handle 32 wavelengths

Hence each fibre can handle 32 traffic signals

so number of fibres needed = 65536/32 = 2048 (D)

9. Traffic doubles every year

So traific after 4 years = 2^4 = 16

And since we need 1 fibre for 1 signal it implies we need = 16 fibres. (C)

10. Similar to question 8

Now we need 16/32 = 0.5

But since fibres must be integers and minimum is 1 hence to transmit 16 signals via a 32 wavelength fibre cable we need 1 fibre. (A)